OpenStudy (yanasidlinskiy):

Solving by the graphing method! Solve the system of equations by graphing. 2x=y+3 4x-3=2y

OpenStudy (yanasidlinskiy):

For the first equation, my solution is: y=(1,-1) y=(0,-3)

OpenStudy (yanasidlinskiy):

For the second one, my solution is: y= (1,1/2) y=(0,-3/2)

OpenStudy (yanasidlinskiy):

What am I doing wrong? Unless I'm misunderstanding how to graph. If that's the case then I'm ready to give up my own brains and use someone elses:)

OpenStudy (anonymous):

Hey.. You will get one solution for both the equations?

OpenStudy (yanasidlinskiy):

It's supposed to be no solution my teacher said.

OpenStudy (anonymous):

Wait..

OpenStudy (yanasidlinskiy):

I'm actually about to blow up and die instead of waiting:)!

OpenStudy (anonymous):

From first equation, bring x and y on one side and similarly do that in second equation, bring constant on right side..

ganeshie8 (ganeshie8):